1.   Zenith Corporation has a 5% $1,000,000 bond issue with a maturity of 12 years. Due to a sudden surge in inflation due to our national debt, the market rate of interest is currently 7%. Interest is paid semi-annually.   Show work!

a.   Will the bond be sold at a “premium”, “par”, or “discount”?
b.   What will be the price and the proceeds?

2.   Four years after the original issue (#1 above), with 8 years remaining on the original bond, all banks are taken over the federal government and the current market rate of interest for all “new” bonds will be 3% (market rate of interest) and the current owner decides to sell the bond and move to Australia which has a stronger form of capitalism.

a.   Will the bond be sold at a “premium”, “par”, or “discount”?
b.   What will be the price and the proceeds?

3.   Eight years after the original issue (#1 above) with 4 years remaining, President Sanders realizes the nationalization of banks by the federal government was not such a good idea, now that China will not lend any more money to the government. Since the country is so much in debt, and everything is free, the market rate of interest is now 12%. Ms. Zhou offers to purchase this bond.

a.   Will Ms. Zhou pay a premium or discount to the original offering and a premium of discount to the secondary offering price?
b.   What will be the price and the proceeds?
Answer #1
1-A Bond will be sold at discount because coupon rate is less than the market interest rate
1-B Value of bond=Using present value function in MS excel pv(rate,nper,pmt,fv,type) rate = 7/2=3.5% nper =12*2 =24 pmt =1000000*5%*1/2 = 25000 fv =1000000 type =0 PV(3.5%,24,-25000,-1000000,0) $839,416.32
price of the bond (839416.32/1000000)*1000 $839.42
proceeds from a bond $839,416.32
2-A Bond will be sold at premium because coupon rate is more than the market interest rate
2-B Value of bond=Using present value function in MS excel pv(rate,nper,pmt,fv,type) rate = 3/2=1.5% nper =8*2 =16 pmt =1000000*5%*1/2 = 25000 fv =1000000 type =0 PV(1.5%,16,-25000,-1000000,0) $1,141,312.64
price of the bond (1141312.64/1000000)*1000 $1,141.31
proceeds from a bond $1,141,312.64
3-A Bond will be sold at discount because coupon rate is less than the market interest rate
3-B Value of bond=Using present value function in MS excel pv(rate,nper,pmt,fv,type) rate = 12/2=6% nper =4*2 =8 pmt =1000000*5%*1/2 = 25000 fv =1000000 type =0 PV(6%,8,-25000,-1000000,0) $782,657.22
price of the bond (782657.22/1000000)*1000 $782.66
proceeds from a bond $782,657.22

AutoSave Off He book 1 - Excel Sign in File Home Insert Page Layout Formulas Data Review View Help O Search Share Comments XAutoSave Off He book 1 - Excel Sign in File Home Insert Page Layout Formulas Data Review View Help O Search Share Comments X